### v1

# Towards Autonomous Agents: Adaptive-planning, Reasoning, and Acting in Language Models

Abstract
We propose a novel in-context learning algorithm for building autonomous decision-making language agents. The language agent continuously attempts to solve the same task by self-correcting each time the task fails. Our selected language agent demonstrates the ability to solve tasks in a text-based game environment. Our results show that the gemma-2-9b-it language model, using our proposed method, can successfully complete two of six tasks that failed in the first attempt. This highlights the effectiveness of our approach in enhancing the problem-solving capabilities of a single language model through self-correction, paving the way for more advanced autonomous agents. The code is publicly available at https://github.com/YenCheHsiao/AutonomousLLMAgentwithAdaptingPlanning.
Index Terms:
In-context learning, large Language Model (LLM), natural language understanding, neural network, pretrained language model.
I Introduction
Large language models (LLMs) are large statistical models that predict the next word, phrase, sentence, or paragraph based on a given input [1]. The quality of the output from a language model can be heavily influenced by the input prompt it receives [2]. One of the capabilities of LLMs is in-context learning, where they learn a new task from a small set of exemplars provided in the prompt during inference [3]. Prompt engineering is the process of designing and refining input prompts to elicit desired responses from LLMs [4].
In Chain-of-Thought (CoT) [5], given a prompt with exemplars that include an input part and an output part, a chain of thought consists of a series of intermediate natural language reasoning steps added between the input and output parts in each exemplar to produce the final output. However, the CoT prompting doesn’t have the ability to update its knowledge from the external world. ReAct [6] prompting address the problem by providing the language model with a prior language description to guide its reasoning about solving diverse language reasoning and decision making tasks and adapting this reasoning by receiving the feedback from the external world. In Reflexion [7], they proposed autonomous decision-making LLM agents by adding a reflection step to the CoT or ReAct prompt to adjust the reasoning, facilitating language agents’ learning from prior failings through verbal reinforcement. VOYAGER [8] is a LLM based agent designed to explore an open-ended world and attain diverse skills through the integration of automatic curriculum, skill library management, and an iterative prompting mechanism incorporating environmental feedback, execution errors, and self-verification to enhance program performance. In Motif [9], an agent is trained through reinforcement learning to maximize rewards from a parameterized model, which is trained based on preferences selected by a language model over pairs of actions, aimed at achieving a specific goal in a given environment.
In this study, we proposed an autonomous decision-making agent with a single language model. Our research contribute to this direction toward autonomy by letting itself to adapt its own policy by correcting its previous failure using its internal knowledge. Our research is different from the work in Reflexion [7], which applied three language models as actor, evaluator, and self-reflection; whereas we propose to use a single LLM to achieve autonomous language agent. Our experiment results show that in fourteen different decision making tasks from the ALFWorld environment, our method can solve the tasks that cannot be completed by the agent using the exemplars from [7] without the reflexion step, demonstrating the effectiveness of our approach.
II Methods
Let a text game be denoted as a function $f$ that maps the state $s\in\mathbb{V}$ and the action $a\in\mathbb{V}$ to an observation $o\in\mathbb{V}$, where $\mathbb{V}$ is a set of vocabulary. Let $\pi_{\theta}$ be an LLM agent over a pre-trained set of parameter $\theta$. Let $s_{0}$ be the initial state of the environment $f$, We aim to produce a sequence of actions $(a_{0},a_{1},a_{2},\dots)$, where $a_{i}\in\mathbb{V}$ for $i\in\mathbb{Z}$, to change the state to a terminal state that indicates the game is cleared.
In ReAct prompting [6], they propose to use an LLM to produce the thought $(t_{0},t_{1},t_{2},\dots)$, where $t_{i}\in\mathbb{V}$ for $i\in\mathbb{Z}$, by $t_{i}\sim\pi_{\theta}(t_{i}|s_{i})$, where $s_{i}=\{s_{i-1},t_{i-1},a_{i-1},o_{i}\}$ for $i\in\mathbb{Z}^{+}$, $a_{i}\sim\pi_{\theta}(a_{i}|s_{i},t_{i})$ for $i\in\mathbb{Z}$, and $o_{i+1}=f(s_{i},a_{i})$ for $i\in\mathbb{Z}$.
The limitation of ReAct prompting [6] is that complex tasks with large action spaces require more demonstrations to learn effectively. The LLM may produce incorrect actions that do not lead to completing the task. Reflexion [7] addresses this problem by using an additional LLM to iteratively provide self-reflection text that will be added to the ReAct prompt for improvement. More specifically, for each trial $ep\in\mathbb{Z}^{+}$, if $ep\geq N$, where $N$ is a maximum number for each tail, a self-reflection $r^{ep}$ is generated and a new state $s^{ep+1}_{0}={s^{ep+1}_{N},r^{ep}}$ is formed to be used as the initial state in the next trial $ep+1$. However, the method in Reflexion [7] necessitate two LLMs, where one LLM is used to generate the thought or action, and another LLM is used to generate the reflection. We will modify this by using a single LLM to generate thought, action, and adaptation, which is the correction from the previous failed trail.
The main idea of our method is for the LLM agent to generate reflections without the need for exemplars. The architecture is shown in Fig. 1. A desire is provided to an agent to motivate it to solve a specific task in a given environment. The agent can perform an action to interact with the environment, causing the state of the environment to change. The agent then receives an observation that describes the status of the environment and a reward signal. The action may be proposed from three different processes: the reasoning process determines the next action based on the current progress; the planning process proposes a series of actions that can be used to solve a specific task; and the adaptation process summarizes previous progress to provide a better plan towards maximizing the reward.
Figure 1: An architecture towards autonomous agent.
We present a novel algorithm in Algorithm 1. Initially, we have the initial state $s_{0}$ which provides instructions, presents exemplars, and describes the environment and the goal for a specific task. $\pi_{\theta}$ is an LLM agent with a set of parameters $\theta$. $\tau=\{s_{0},t_{0},a_{0},o_{1},\dots\}$ is a sequence of the concatenation of state, thought, action, and observation, where $s_{k}$, $t_{k}$, $a_{k}$, and $o_{k}$ are sequences of tokens representing the $k$-th state, thought, action, and observation for $k\in\mathbb{Z}$, respectively. The return $R(\tau)$ is a string indicating whether the task is completed or not. $ep$ is a variable indicating the number of trials. The environment is reinitialized at each trial.
At the first time step $k=0$, the thought is sampled from
$t^{1}_{0}\sim\pi_{\theta}(t^{1}_{0}|s^{1}_{0}),$ (1)
where the subscript 00 indicates the first time step, and the superscript $1$ indicates the first trail. $t^{1}_{0}$ represents the first thought in the first trial, and $s^{1}_{0}$ represents the first state in the first trial, with both being sequences of tokens. The action is then sampled from
$a^{1}_{0}\sim\pi_{\theta}(a^{1}_{0}|s^{1}_{0},t^{1}_{0}),$ (2)
where $a^{1}_{0}$ represents the first action in the first trial, and the action is a sequence of tokens. The second observation in the first trial, $o^{1}_{1}$, is a sequence of tokens obtained by executing the action $a^{1}_{0}$ in the environment $f$ at state $s^{1}_{0}$ as
$o^{1}_{1}=f(s^{1}_{0},a^{1}_{0}).$ (3)
A new state $s^{1}_{1}$ is formed by concatenating the thought $t^{1}_{0}$, action $a^{1}_{0}$, and observation $o^{1}_{1}$ after state $s^{1}_{0}$ as
$s^{1}_{1}=\{s^{1}_{0},t^{1}_{0},a^{1}_{0},o^{1}_{1}\}.$ (4)
If a maximum time step is reached, the task fails and the return $R(\tau)$ is concatenated with ”New plan: ”. They are concatenated after the current state of the environment $s^{1}_{49}$ to form the initial state in the next trial $s^{2}_{0}$ as
$s^{2}_{0}=\{s^{1}_{50},R(\tau)\},$ (5)
where $\tau=\{s_{0},t_{0},a_{0},o_{1},t_{1},a_{1},o_{2},t_{2},\dots o_{50}\}$. In the next trial, the first thought in the second trial, $t^{2}_{0}$, is sampled from the LLM by
$t^{2}_{0}\sim\pi_{\theta}(t^{2}_{0}|s^{2}_{0}),$ (6)
We call the initial thought $t^{ep}_{0}$ at the $ep$-th trial for $ep>1$ as the adaptation from the ($ep-1$)-th trail and $t^{ep}_{0}$ indicates the correction of the ($ep-1$)-th failed trail to improve the next trail. In the next step, we propose to replace the initial state in the second trial with the initial state in the first trail to reduce the context length. We call this step compression. By performing compression, the first action in the second trail will only be conditioned on the initial state in the first trail $s^{1}_{0}$ and the adaptation from the first trail $t^{2}_{0}$ as
$a^{2}_{0}\sim\pi_{\theta}(a^{2}_{0}|s^{2}_{0},t^{2}_{0}).$ (7)
Algorithm 1 Reflexion [7] with one language model.
  Initialize the world state $s_{0}$ as a text of exemplars and task, where each token $\in Vocab$.
  Let $\pi_{\theta}$ be a LLM agent over a pre-trained set of parameters $\theta$.
  Let a trajectory $\tau=\{s_{0},t_{0},a_{0},o_{1},\dots\}$ be a sequence of state, thought, action, and observation.
  Let $R(\tau)$ be the return for trajectory $\tau$.
  Let $ep=1$.
  While $R(\tau)\neq"OK"$ do
        Let $k=0$.
        While $k<50$ ||||| | $R(\tau)="OK"$ do
            Generate thought $t^{ep}_{k}\sim\pi_{\theta}(t^{ep}_{k}|s^{ep}_{k})$.
            Compression step:
                If $k=0$, then $s^{ep}_{0}=s_{0}$.
            Generate action $a^{ep}_{k}\sim\pi_{\theta}(a^{ep}_{k}|s^{ep}_{k},t^{ep}_{k})$.
            Get observation $o^{ep}_{k+1}=f(s^{ep}_{k},a^{ep}_{k})$.
            Let $s^{ep}_{k+1}=\{s^{ep}_{k},t^{ep}_{k},a^{ep}_{k},o^{ep}_{k+1}\}$.
            $k:=k+1$
        Concatenate $R(\tau)$ with ”New plan: ”.
        $s^{ep+1}_{0}=\{s^{ep}_{k},R(\tau)\}$
        $ep:=ep+1$
III The ALFWorld environment
There are six types of tasks in the ALFWorld environment [10]: Pick and Place, Examine in Light, Clean and Place, Heat and Place, Cool and Place, and Pick Two and Place. For each task, a description of available receptacles is given in the first part of the instruction as follows: You are in the middle of a room. Looking quickly around you, you see a {recep1 id}, a {recep2 id}, …, and a {recepN id}, where recepN refers to the Nth receptacles like drawers or cabinet and $\in\mathbb{Z}^{+}$. An example is shown in the text in Fig. 2 with a green background.
After the description of the available receptacles, the goal instructions are provided based on the six different task types. For a Pick and Place task, the instruction will be either ”put a obj in recap” or ”put some obj on recap”. For an Examine in Light task, it will be either ”look at obj under the lamp” or ”examine the obj with the lamp”. For a Clean and Place task, it will be either ”put a clean obj in recap” or ”clean some obj and put it in recap”. For a Heat and Place task, it will be either ”put a hot obj in recap” or ”heat some obj and put it in recap”. For a Cool and Place task, it will be either ”put a cool obj in recap” or ”cool some obj and put it in recap”. For a Pick Two and Place task, it will be either ”put two obj in recap” or ”find two obj and put them in recap”. Inside the curly brackets, {obj}, {recep} and {lamp} refer to object, receptacle, and lamp classes, respectively. An example is shown in the text in Fig 2 with a red background.
After the goal instruction of the task, the agent or the user can interact with the game environment using the following nine different text actions: go to recap id, open recap id, clean obj id with recap id, take obj id from recap id, close recap id, close recap id, heat obj id with recap id, put obj id in/on recap id, and use recap id. Given the action, the game environment will return text observations accordingly. An example is shown in the text in Fig. 2 with a magenta background.
Figure 2: A trajectory in the ALFWorld environment [10]. The text in the black box is composed of the description of available receptacles, the description of the goal instruction, and a sequence of actions and observations. (1) The description of available receptacles is listed in the first part of the text with a green background. (2) The description of the goal instruction with a red background shows that the task is a Cool and Place task, and the goal is to cool some pan and put it in countertop. (3) The sequence of actions and observations with a magenta background shows the actions performed and the corresponding observations from the environment. The actions are in boldface after the greater than symbol, and the observations are in regular text below each action.
IV Experimental designs and results
IV-A Experiment on ReAct prompting in the Alfworld environment
In ReAct [6], they randomly annotate the trajectories for each task type. Each trajectory includes sparse thoughts that decompose the goal, track subgoal completion, determine the next subgoal, and use commonsense reasoning to find an object and determine what to do with it. An exemplar used in ReAct [6] is shown in Fig. 3. The instruction text, “Interact with a household to solve a task. Here are two examples.”, is added above the exemplars to indicate that the general goal is to complete a household task, and the texts below it include two exemplars. Following the examples, the text ’Here is the task.’ is concatenated, followed by the description of available receptacles and the description of the goal instruction from the ALFWorld environment [10]. Finally, the greater-than symbol (’$>$’) is added to indicate that the language model should generate text after it, starting with actions or thoughts.
Figure 3: An exemplar in ReAct [6] for a Heat and Place task in the ALFWorld environment [10]. The text in the black box is composed of the description from the ALFWorld environment [10] and the thoughts annotated by ReAct [6]. The thought that decomposes the goal is shown with a magenta background. The thought that uses commonsense reasoning to find an object and determine what to do with it is shown with a red background. The thoughts that track subgoal completion are shown with a cyan background. The thoughts that determine the next subgoal are shown with a green background.
ReAct [6] evaluated their method on 134 tasks in the ALFWorld environment [10], achieving a success rate of 70.9% using PaLM-540B [11] and 78.4% using the GPT-3 text-davinci-002 model [12], where each output token is selected with the highest probability. However, PaLM-540B [11] is not publicly available, and the text-davinci-002 model [12] was shut down by OpenAI on January 4th, 2024. Due to the difficulty in reproducing the results from [6], we tested various open-source LLMs and selected the one with the highest success rate to develop our method. The results in Table I show that the gemma-2-9b-it model outperforms other models with a success rate of 62% in a 12-hour run. We will use gemma-2-9b-it for the subsequent experiments.
TABLE I: Success rate of different open-source language models with ReAct [6] applied to the ALFWorld environment [10] in a 12-hour run for each model
Developer Name of the language model Success rate (%) Number of success tasks number of tasks
Google gemma-2-9b 40 25 62
Google gemma-2-9b-it 62 37 59
Mistral AI Mistral-7B-v0.3 28 16 57
Mistral AI Mistral-7B-Instruct-v0.3 25 14 54
LLaMA Llama-2-7b-hf 8 4 50
Microsoft Phi-3-medium-128k-instruct 41 12 29
DeepSeek deepseek-llm-7b-base 14 15 103
Hugging Face H4 zephyr-7b-alpha 15 8 52
IV-B Experimental on Reflexion with one LLM
In Reflexion [7], a status text indicating whether a task is successful or not, along with a reflection text that guides the next trial toward success, are concatenated after the ReAct exemplars. Reflexion [7] used two exemplars, as shown in Fig. 4, to guide the LLM in generating the reflection.
Figure 4: Two exemplars in Reflexion [7] for the ALFWorld environment [10]. The text in each black box comprises one exemplar from Reflexion [7] designed to guide an LLM in generating the correct action to complete a task in the ALFWorld environment [10]. In each black box, the text preceding the yellow-background text represents a ReAct trajectory, as shown in Fig. 3. The text next to ”STATUS: ” indicates whether the task is completed. If the task is completed, the yellow-background text will read ”STATUS: OK”. If the task is not completed, it will read ”STATUS: FAIL”. The reflection text is highlighted with a cyan background.
Unlike Reflexion [7], which using one LLM to generate thoughts and actions and another LLM to generate reflections, we employ a single LLM to generate thoughts, actions, and reflections by concatenating the two Reflexion exemplars [7] after the two ReAct exemplars [6] for each task. This approach allows the reflection part to be generated after the LLM receives the string ”STATUS: FAIL”. We experiment with this single-LLM setup using the exemplars from ReAct [6] and Reflexion [7] as depicted in Fig. 5.
Initially, the input to the LLM is a prompt with the instruction, exemplars, and the description of the environment and the task to be completed. The output of the LLM will be the input prompt concatenated with a series of thoughts, actions, and observations. We extract the first thought or action from the output by extracting the line immediately following the input prompt content. The line immediately after the greater than symbol ($>$) is used as input to the environment. If the input to the environment contains the string ”think:”, we overwrite the output with ”OK.”, following the process described [6, 7].
If the task is completed, the environment will return a variable with a value of $1$, and the process is finished. Otherwise, we concatenate the extracted thought or action with the observation and append this new text to the input prompt forms the next input prompt. After $49$ steps, if the task remains unfinished, we reinitialize the ALFWorld environment and append ”STATUS: FAIL” and ”New plan:” after the last observation. This updated text becomes the next input prompt to the LLM. The output of the LLM will include the reflection text following ”New plan:”, which is extracted and concatenated above the initial prompt, $s^{ep}_{0}$. This new prompt is then used as the next input to the LLM to obtain corrected thoughts and actions. We set a maximum of $10$ trails; if $ep\geq 10$, we terminate the trial and proceed to the next task. The result of this experiment, conducted over a 24-hour run using gemma-2-9b-it, are shown in the second column of Table II.
Figure 5: The process of using a single LLM to generate thoughts, actions, and reflections by applying ReAct’s exemplers [6] and Reflexion’s exemplars [7] for solving tasks in the ALFWorld environment [10].
TABLE II: Number of steps for different methods to complete tasks in the ALFWorld environment [10]
Task number One LLM with exemplars from ReAct [6] and Reflexion [7]
1 13
2 12
3 fail
4 10
5 17
6 87
7 fail
8 19
9 fail
10 10
11 17
12 67
13 fail
14 10
V Discussion
In Section IV-A, we observed that gemma-2-9b-it model outperforms other models (gemma-2-9b, Mistral-7B-v0.3, Mistral-7B-Instruct-v0.3, Llama-2-7b-hf, Phi-3-medium-128k-instruct, deepseek-llm-7b-base, and zephyr-7b-alpha) in solving various tasks in the ALFWorld environment [10] using ReAct prompting [6]. We identified three common issues with these LLMs. First, the LLMs may attempt to retrieve an object from a location where the object does not exist, repeatedly performing the same action until reaching the maximum number of steps, $49$. Second, the LLMs may select an incorrect item. For instance, when the task is to ”put a clean cloth in countertop”, the gemma-2-9b model may pick up a ”handtowel” instead. Subsequently, the LLM cannot complete the task by cleaning the ”handtowel”, resulting in repeated attempts to clean and place the ”handtowel” in the ”countertop” until the maximum step limit is reached. Third, some LLMs will misinterpret the order of sub-goals. For example, when the goal is to ”examine the cd with the desklamp”, the gemma-2-9b-it model erroneously attempts to retrieve a ”desklamp” first. After failing to obtain the ”desklamp”, the LLM searches for the ”cd” but only revisits previously searched locations instead of exploring new ones.
In Section IV-B, we investigated the use of a single gemma-2-9b-it model to generate actions, thoughts, and reflections by incorporating exemplars from ReAct [6] and Reflexion [7] into the input prompt. The experiment was conducted over 24 hours and involved six different tasks. Among the six tasks, five were completed successfully, while one were not finished after reaching 10 reflection trials. For the five tasks that were completed, there were four tasks finished before reaching the reflection step.
In the failed task (task number 3), we only reflection obtained was: ”I was stuck in a loop in which I continually examined the fridge 1 instead of using a different action. I should have looked for a lettuce in the fridge 1, then taken it. I will try to execute a different action if I am stuck in a loop again.” for every trial. However, the LLM did not attempt different actions in subsequent trials and continued to examine the fridge. The actual issue in task 3 is that the fridge does not contain lettuce, but the LLM incorrectly assumes it is present and persistently attempts to retrieve it.
In the sixth task (task number 6), the task could not be completed in the first trial but was successfully completed in the second trial with 38 steps. The goal of this task is to ”find two pillow and put them in sofa”. In the first trial, the agent failed to complete the task because it attempted to pick up ’pillow 2’ from ’sofa 1’; however, ’sofa 1’ only contained ’pillow 1’. Afterward, the agent continued trying to put ’pillow 2’ on ’sofa 1,’ but it failed because it did not have the pillow. Subsequently, the LLM did not output any text until the end of the trial. After this, the following adaptation was generated and appended to the input prompt for the second trial: ”I was stuck in a loop in which I continually looked for the second pillow in sofa 1. I should have looked for the second pillow in armchair 1, sidetable 1, and cabinet 1-4. I will try to execute a different action if I am stuck in a loop again.”. In the second trial, the agent found that ’pillow 2’ was on ’armchair 1,’ picked it up, placed it on the sofa, and completed the task. The trajectories of the two trials in task number 6 are shown in Fig. 6.
Figure 6: Two trajectories in task number 6 are shown in Table II. The text in each black box represents one trajectory in the ALFWorld environment [10], as detailed in the experiment in Section IV-B. In each black box, the text at the top includes the description of available receptacles and the goal instruction, as mentioned in Fig. 2. Below the goal instruction, the text following the greater-than symbol (’$>$’) represents the thoughts or actions generated by the gemma-2-9b-it LLM. Text beginning with ’$>$ think:’ indicates a thought, while all other text beginning with without ’think’ represents actions. The text beneath each thought or action is the observation from the ALFWorld environment [10]. The trajectory of the first trial in task number 6 is shown in the left black box. The LLM agent failed to complete the task because it couldn’t find the second pillow, incorrectly believing it had finished the task, and then continued generating empty strings. The adaptation the LLM agent generated is highlighted with a cyan background. In the second trial, shown in the right black box, the adaptation is appended above the description of the available receptacles. The agent successfully found the second pillow in the second trial and completed the task.
In the twelfth task (task number 12), the task could not be completed in the first trial but was successfully completed in the second trial with 18 steps. The goal of this task is to ”put a cool tomato in microwave”. In the first trial, the agent failed to complete the task because it attempted to take a ’tomato’ from ’fridge 1’; however, ’fridge 1’ doesn’t contained any tomato. Afterward, the agent continued trying to take a ’tomato’ from ’fridge 1’ until the end of the trial. After this, the following adaptation was generated and appended to the input prompt for the second trial: ”I was stuck in a loop in which I continually looked for a tomato in the fridge. I should have looked for a tomato in a different environment. I will try to look for a tomato in a different environment in the next trial.”. In the second trial, the agent found that ’tomato 1’ was on ’countertop 1,’ picked it up, cooled it with the fridge, placed it in the microwave, and completed the task. The trajectories of the two trials in task number 12 are shown in Fig. 7.
Figure 7: Two trajectories in task number 12 are shown in Table II. The text in each black box represents one trajectory in the ALFWorld environment [10], as detailed in the experiment in Section IV-B. In each black box, the text at the top includes the description of available receptacles and the goal instruction, as mentioned in Fig. 2. Below the goal instruction, the text following the greater-than symbol (’$>$’) represents the thoughts or actions generated by the gemma-2-9b-it LLM. Text beginning with ’$>$ think:’ indicates a thought, while all other text beginning with without ’think’ represents actions. The text beneath each thought or action is the observation from the ALFWorld environment [10]. The trajectory of the first trial in task number 6 is shown in the left black box. The LLM agent failed to complete the task because it couldn’t find a tomato, and then continued checking the same receptacle. The adaptation the LLM agent generated is highlighted with a cyan background. In the second trial, shown in the right black box, the adaptation is appended above the description of the available receptacles. The agent successfully found a tomato in the second trial and completed the task.
VI Conclusion
We present a novel in-context learning algorithm designed for a single language model to complete tasks in a text-based game by correcting its previous failures. This approach reduces the number of models used in previous work [7] from two LLMs to one LLM. Our findings indicate that the gemma-2-9b-it model achieves the highest success rate of 62% for completing tasks in the ALFWorld environment [10] using ReAct prompting [6], compared to other selected open-source language models. We show that using gemma-2-9b-it, two of the six tasks that could not be completed in one trial can be completed in the second attempt by appending the adaptation from the previous trial. Future work will involve further experimentation with different approaches to completing decision-making tasks using a single language model without Reflexion exemplars to reduce the number of tokens in the input prompt.
Acknowledgments
The computational work for this project was conducted using resources provided by the Storrs High-Performance Computing (HPC) cluster. We extend our gratitude to the UConn Storrs HPC and its team for their resources and support, which aided in achieving these results.